OpenStudy (anonymous):

Calculate the median of the following set of data: 67, 38, 28, 39, 94, 22, 19, 30, 84, 64

OpenStudy (uri):

What's the middle number?

OpenStudy (uri):

@xo_ellaaavasquez

OpenStudy (anonymous):

put all the numbers in order from smallest to largest

OpenStudy (anonymous):

the middle ones are 38 and 39

OpenStudy (anonymous):

you have to list the numbers in order from least to greatest

OpenStudy (anonymous):

i did but i got it wrong do i have to add up 38 and 39 or what do i do

OpenStudy (anonymous):

with the 2 remaining numbers in the middle, you have to add them together and then divide by 2 which gives you 77/2= 38.5*

OpenStudy (anonymous):

38.5

OpenStudy (uri):

add 38 and 39 and divide by 2

OpenStudy (anonymous):

ohhh i see i knew i was suppose to divide but i didnt know it was by 2 thanks ya'll

OpenStudy (uri):

38+39=77/2 =38.5